# Mahyar Habibi,1 Dirk Hovy,2 Carlo Schwarz1

### v1

@@ 3.3 An Alternative Approach to Content Moderation
We propose an alternative approach to tackle the issue of online toxicity that has the potential to mitigate some of the distortions in online discourse. Instead of removing toxic content outright, content moderators could instead use the language generation capabilities of large language models and rephrase the message of Tweets using less toxic language. This makes it possible to reduce the toxicity of online content while keeping the original content as far as possible intact. [^17] The detoxification of online content has recently also become a task tackled by computer scientists in the TextDetox 2024 competition.
@@
Note that we do not argue that all content should necessarily remain on the platform. There might well be content and thoughts that should not be admitted to be shared on online platforms. Our argument rather is that there exist significant amounts of content that violate the content standards of social media platforms (e.g., grave insults) but, at the same time, might contain political expression or information that might be worth preserving. In such cases, rephrasing might be a suitable approach. For example, the rephrasing of messages has been shown to facilitate communication in partisan politics (Argyle et al., 2023).
@@
We test the potential of such an approach using GPT4o-mini to rephrase Tweets with a toxicity score above 0.5. We provide additional details on the rephrasing prompt in Section B.2. We also show examples of rephrased Tweets in Table A.3. Overall, GPT4o-mini can rephrase Tweets in a significantly more civil manner while keeping the main message intact. Rephrasing reduces the toxicity of Tweets from 0.71 to 0.26.
@@
We then compare the BCD for two content moderation strategies. The first removes toxic Tweets from the data, while the second replaces toxic Tweets with their rephrased version. The results from this analysis are shown in Figure 5. We find that in contrast to removal, the rephrasing of Tweets leads to smaller changes in the BCD. The fact that the gap is widening with lower content-moderation thresholds strikes us as intuitive as rephrasing highly toxic content that only contains insults leads to similar distortions than outright removal. Only once we consider content that contains other content independent of the toxic language, rephrasing achieves its intended goal of preserving the original content. This suggests that such an approach can reduce the toxicity of online content with fewer distortions to online content.
@@ 4 Conclusion
This paper proposes and validates a new methodology to measure content-moderation-induced distortions to online content. This new methodology enables us to, for the first time, document the impact of the removal of toxic content on online discourse. Given the crucial importance and heated nature of the debate surrounding this issue, it is important to be clear about what our measure achieves and what it does not capture. The measure captures the distortions in the embedding space of the content that is circulating on online platforms. As the embeddings measure many semantic characteristics of a text, changes in the mean and the variance of the embeddings provide insight into the extent of the semantic distortions.
That being said, our measure does not, nor is it the goal of this paper, provide a universal measure of online content and free speech. Given the complex and multifaceted nature of the concept of free speech that encompasses different philosophical and legal standpoints (see for example Warburton, 2009, for a review), no one measure could ever capture all facets of online discourse. In the same way, despite extensive efforts by the research community, no measure of hate speech could ever capture the importance of cultural context and changing societal norms (e.g., Brown, 2017).
Nonetheless, automated hate speech detection tools have proven useful to rain in toxic speech and are hence widely deployed online. As online platforms and regulators inevitably face trade-offs when it comes to the moderation of online content, we believe it is important to have measures that are also able to capture the cost of content moderation. We believe that by shedding light on the trade-offs involved in content moderation, our measure provides a fundamental advancement in the application of NLP tools for the problem of content moderation and highlights a highly fruitful direction for future research.
@@ Appendix A Additional Details on Data
In our study, we initiated our data collection process with a cohort of 432,882 randomly selected American Twitter users, a dataset that was meticulously collected in 2015 by (Siegel et al., 2021). This particular sample selection bears two pivotal advantages. Firstly, it affords us the opportunity to construct a comprehensive and representative overview of Twitter activity across the United States. Secondly, these users have been actively engaged on Twitter for several years, and we do not face problems with composition changes. Consequently, we were able to conduct an insightful analysis based on a good approximation of the content that circulated on Twitter. The initial step in our data collection involved retrieving the Tweets posted by each user within the (Siegel et al., 2021) sample. In totality, our dataset encompassed 399 Million Tweets spanning from 2014 to the commencement of 2022, along with corresponding user profile information. We provide some summary statistics in Table A.1
@@
We developed a political Tweet classifier by fine-tuning the BERTweet model on a human-labeled dataset of political and non-political Tweets. To create the training dataset, we first compiled a list of political keywords and accounts associated with political figures. This produced a subset of data with a higher density of political Tweets, though many non-political Tweets remained. We then randomly sampled 10,000 Tweets and distributed them among five undergraduate research assistants for manual labeling. Coders were provided with labeling criteria and instructed to categorize each Tweet in a binary true/false format. Each subset included 200 shared Tweets to evaluate inter-coder reliability. We evaluated agreement using Cohen’s kappa coefficient, which ranges from 0 to 1, where 1 indicates perfect agreement and values near 0 suggest agreement no better than chance. The average Cohen’s kappa score across all coder pairs was 0.76. We then use this dataset to train and evaluate our political Tweet classifier.
@@
To gauge the toxicity of these Tweets, we use Google’s Perspective API, a tool widely acknowledged for its efficacy in identifying hate speech (Wulczyn et al., 2017; Dixon et al., 2018). This state-of-the-art API assigns a toxicity score ranging from 0 to 1 across six distinct dimensions: general toxicity, severe toxicity, identity attacks, insults, profanity, and threats. The scores can be approximately interpreted as the probability of a randomly chosen user classifying content as toxic. For example, a score of 0.8 means that around 80% of users would judge the content to be toxic.
The Perspective API performs well in classifying toxic text and can assess Tweets in several languages, including English, Spanish, French, German, Portuguese, Italian, and Russian. In our dataset, English, unsurprisingly, is the predominant language, and we restrict our analysis to English-speaking Tweets. Instances where the API did not assign toxicity scores were primarily due to the absence of textual content, such as Tweets containing only hyperlinks. We provide examples of highly toxic Tweets in Table A.3. The examples indicate that the Perspective API accurately identifies toxic content. As described in the main paper, we additionally use the Detoxify (Hanu and Unitary team, 2020) package and OpenAI’s Moderation API (OpenAI, 2024) as an alternative toxicity classification algorithm. Overall, we also find that the different models broadly agree in their toxicity evaluation of Tweets. The correlation between toxicity scores from the Perspective API and Detoxify is 0.85, while the correlation between the Perspective API and the Moderation API is 0.72.181818As the Moderation API does not report a direct Toxicity score, we always consider the maximum of all provided scores.
@@
The Top2Vec algorithm proceeds in three steps. First, the texts are transformed into embedding vectors based on pre-trained models. In our particular case, we make use of the universal-sentence-encoder model (Cer et al., 2018). Second, the embeddings are projected into lower dimensional space using UMAP (McInnes et al., 2018). This helps to overcome the sparsity of the higher dimensional space. Third, the reduced embeddings are clustered into topics using HDBSCAN (McInnes et al., 2017), which identifies the centroids for each topic vector. We specify that HDBSCAN should only create clusters with at least 1500 observations. Top2Vec creates a unique assignment between texts and topics. The topic words are then derived based on the word vectors that are closest to the topic centroid. We report the most important topics words as well as well as the assigned topic labels in the table below.
@@ B.1. Additional Details on Embeddings
To distill information from the raw text of Tweets into quantifiable data, we use three pre-trained models: BERTweet (Nguyen et al., 2020), RoBERTa (Liu et al., 2019), and DeBERTa (He et al., 2021). These models all built on the BERT-style Transformer architecture, which has become a staple in text classification and natural language understanding tasks. The strength of these models comes from their utilization of an ”attention” mechanism, allowing them to evaluate words in the context of their surrounding text, thereby capturing the subtleties of language that are often lost in traditional analysis. We transform the raw Tweets into analyzable embeddings by tokenizing the text into its constituent word tokens. These tokens then serve as input to our models which produce numerical representations—or embeddings—of each Tweet. These embeddings convey the semantic and syntactic nuances of the language used by Twitter users and form the backbone of our computational analysis. Further details on the specific attributes of these models within our study are provided in the following.
@@
The third model we are using in our analysis is DeBERTa (He et al., 2021). DeBERTa (Decoding-enhanced BERT with disentangled attention) introduced innovative mechanisms that refine the workings of BERT and RoBERTa models. Its distinctive feature lies in the disentangled attention mechanism, which considers the content and the position of words separately, offering a more nuanced understanding of the text. Each word is represented by dual vectors that capture what the word is and where it stands in a sentence. This allows for a better interpretation of language nuances. Furthermore, DeBERTa’s enhanced mask decoder predicts masked tokens by utilizing their absolute positions, an improvement that aids in the pre-training process. With these advancements, DeBERTa improved the model’s pre-training efficiency and improved the performance across a variety of downstream tasks, like the generation of new text that mimics human speech.
@@ Appendix C Additional Results
The following section describes additional results and robustness checks. Specifically, we benchmark the content-moderation-induced distortions relative to the maximum possible distortion that can be created when removing a specific number of Tweets. Further, we repeated our main analysis using alternative 1) embedding models, 2) toxicity scores, and 3) samples. We also show that our results are not driven by Tweets belonging to the “Insult” topic as created by the Top2Vec topic model.
Benchmarking BCD
To benchmark the magnitude of the content-moderation-induced distortions, we compare changes in the BCD when removing Tweets based on their toxicity scores versus removing an equal number of Tweets with the greatest Euclidean distance from the centroid of all embedded Tweets. Given the definition of Bhattacharyya distance, removing Tweets furthest from the distribution’s center approximates the upper bound of BCD changes possible by removing a specific number of Tweets. Figure Figure C.1 shows the changes in BCD when removing Tweets with toxicity scores above a certain threshold and compares this to removing an equal number of Tweets based on their distance from the centroid. The results indicate that removing toxic Tweets increases BCD by approximately 20% of the maximum increase from removing Tweets with a large Euclidean distance.
Figure C.1: Benchmarking BCD
@@
First, we reproduce our findings using the alternative embeddings described in Section B.1. This test rules out that our findings are driven by the particularities of the specific transformer model we have chosen, even though BERTweet is one of the standard choices for the analysis of English-speaking Twitter data. For this robustness exercise, we created new embeddings based on the RoBERTa and DeBERTa models and reconstructed the BCD based on these embeddings. The findings in Figure C.2 highlight that the findings are remarkably similar not only with regards to the overall patterns but also the magnitudes of the content-moderation-induced increases in the BCD.
@@
Notes: The figure shows the BCD computed using embeddings generated by DeBerta and RoBerta, after the exclusion of toxic and random Tweets from the data.
@@
As a last robustness check, we repeat our main analysis based on three alternative samples of Tweets. First, we use 1 million randomly drawn Tweets from our English data without filtering for political content. Second, we use 1 million randomly drawn Tweets from a sample of politically interested German Twitter users. Third, we use 1 million randomly drawn Tweets from a sample of politically-interested Italian Twitter users. To create embeddings for the German and Italian samples, we use the “bert-base-german-uncased” and “twitter-xlm-roberta-base” models, respectively. Toxicity in all cases was coded using the Perspectives API. For each of these samples, we repeat the analysis from Figure 1. The results are presented in Figure C.4. We find that independent sample content moderation appears to introduce distortions to the semantic space as measured by the BCD.
@@
The figure below reproduces Figure 1 when we remove all Tweets belonging to the “Insult” topic in advance. The results are virtually identical, suggesting again that content moderation also introduces distortions of the semantic space over and above purely removing uncivil content.

### v2

@@ 3.2 What Drives Distortions to Online Content
In the first analysis, we show that it is possible to reduce the toxicity of online content while creating smaller distortions to the semantic space. We thereby highlight that there is content that can be salvaged from toxic Tweets. For this analysis, we propose an alternative approach to address the issue of online toxicity, which has the potential to mitigate some of the distortions in online discourse. Instead of removing toxic content outright, content moderators could use the language generation capabilities of LLMs to rephrase the message of Tweets using less toxic language. This transformation reduces the toxicity of online content while preserving the original content as much as possible. [^16] The detoxification of online content has recently also become a task tackled by computer scientists in the TextDetox 2024 competition.
@@
Note that we do not argue that all content should necessarily remain on the platform. There might well be content and thoughts that should not be admitted to be shared on online platforms. Some posts are beyond salvaging. Our argument, rather, is that a substantial share of content violating platform standards (e.g., grave insults) or toxicity thresholds may nonetheless carry political expression or information worth preserving. In such cases, rephrasing might be a suitable approach. For example, the rephrasing of messages has been shown to facilitate communication in partisan politics (Argyle et al., 2023).
@@
We test the potential of such an approach using GPT4o-mini to rephrase Tweets with a toxicity score above 0.5. We provide additional details on the rephrasing prompt in Section B.2. We also show examples of rephrased Tweets in Table 1. Overall, GPT4o-mini can rephrase Tweets in a significantly more civil manner while retaining the main message. Rephrasing reduces the toxicity of Tweets from 0.71 to 0.26, while maintaining very similar content (i.e., the average cosine similarity between rephrased Tweets and the originals is 0.97).
@@
We then compare the BCD for two content moderation strategies. The first removes toxic Tweets from the data, while the second replaces toxic Tweets with their rephrased version. The results from this analysis are shown in Figure 5. We find that, in contrast to removal, the rephrasing of Tweets leads to smaller changes in the BCD. If the content changes were solely driven by the presence of toxic language, replacing Tweets with a rephrased version should cause the same change to the semantic space. However, since we find that rephrasing can mitigate some of the distortions, it suggests that the observed distortions are not merely artifacts of toxic language itself.
The fact that the gap is widening with lower content-moderation thresholds strikes us as intuitive, as rephrasing highly toxic content that only contains insults leads to similar distortions as outright removal. These are the posts that cannot be salvaged. Only once we consider Tweets that contain other content, independent of the toxic language, does rephrasing achieve its intended goal of preserving the original content. This suggests that rephrasing can reduce the toxicity of online content with fewer distortions. Note that we would not expect the BCD of the rephrased Tweets to be zero, as rephrasing introduces changes in the language of Tweets.
@@
Accounting for the Toxicity Dimension of the Embedding Space
As a second analysis, we directly remove the toxicity component from the embedding space. To do so, we build on the literature on the debiasing of embeddings (e.g., Bolukbasi et al., 2016; Liang et al., 2020) and create orthogonal projections of the embeddings matrix $\mathbf{X}$ with respect to the Toxicity scores. In other words, we remove any variation in the embeddings of Tweets that can be explained by their toxicity. Any remaining variances in the embeddings should, therefore, capture differences in the content of the Tweets, independent of their toxicity.
Initially, we construct a “toxicity dimension” by calculating the average vector differences between the embeddings of highly toxic Tweets and their rephrased counterparts. Next, we use Principal Component Analysis (PCA) to identify the 10 principal vectors that characterize toxicity. The final step involves projecting the embedding space so that it is orthogonal with respect to the toxicity subspace defined by the principal components. This procedure produces an orthogonalized matrix $\mathbf{\tilde{X}}$, which is uncorrelated with the toxicity scores. In Online Appendix C, we also present an alternative method to remove the toxicity dimension from the embedding space based on regression residuals, which yields similar results.
We then repeat the previous analysis using the orthogonalized embedding matrix $\mathbf{\tilde{X}}$. Figure 6 compares the BCD obtained when removing Tweets from the original embedding space (orange line) and the orthogonalized embedding space (blue line). We also report the BCD for the removal of rephrased Tweets (red line). [^17] In this exercise, all Tweets are first replaced with their rephrased versions to define the baseline sample. Afterwards, Tweets are removed based on their pre-rephrasing toxicity scores.
Three key findings emerge from this analysis. First, even in the orthogonalized embedding space, we continue to observe substantial, albeit somewhat smaller, changes in the BCD. This again suggests that the observed shifts in BCD arise to a considerable extent from the underlying content of Tweets rather than their toxic language. Second, the BCD resulting from the removal of orthogonalized Tweets is virtually identical to that obtained by removing their rephrased counterparts. This demonstrates that our orthogonalization procedure effectively removed the toxicity dimension, bringing the embeddings close to those of the rephrased Tweets. Third, removing rephrased Tweets produces more pronounced changes to the embedding space than replacing toxic Tweets with their rephrased versions (see Figure 5). Intuitively, this underscores that removing content introduces larger distortions to the embedding space than rephrasing.
Figure 6: Controlling for Toxicity
Notes: The figure shows the BCD, derived from the original embedding space (orange line) and the orthogonalized embedding space (blue line). We also report the BCD for the removal of rephrased Tweets (red line).
@@ 4 Conclusion
This paper proposes and validates a new methodology for measuring content-moderation-induced distortions in online content. This new methodology enables us, for the first time, to document the impact of removing toxic content on online discourse. Given the crucial importance and heated nature of the debate surrounding this issue, it is important to be clear about what our measure achieves and what it does not capture. The BCD provides a heuristic measure for changes in the embedding space of content circulating on online platforms. As the embeddings measure many semantic characteristics of a text, changes in the mean and the variance of the embeddings provide insight into the extent of the semantic distortions.
That being said, our measure does not provide a universal measure of online content and free speech (nor is it the goal of this paper). Given the complex and multifaceted nature of the concept of free speech, which encompasses different philosophical and legal standpoints (see, for example, Warburton, 2009, for a review), no single measure can ever capture all facets of online discourse. Similarly, despite extensive efforts by the research community, no measure of hate speech can ever fully capture the importance of cultural context and changing societal norms (e.g., Brown, 2017).
Nonetheless, automated hate speech detection tools have proven helpful to rein in toxic speech and are hence widely deployed online. As online platforms and regulators inevitably face trade-offs when it comes to moderating online content, we believe it is essential to have measures that also capture the cost of content moderation. We believe that by shedding light on the trade-offs involved in content moderation, our measure represents a fundamental advancement in the application of NLP tools for content moderation and highlights a highly fruitful direction for future research.
@@ Appendix A Additional Details on Data
In our study, we initiated our data collection process with a cohort of 432,882 randomly selected American Twitter users, a dataset that was meticulously collected in 2015 by (Siegel et al., 2021). This particular sample selection bears two pivotal advantages. Firstly, it allows us to construct a comprehensive and representative overview of Twitter activity across the United States. Secondly, these users have been actively engaged on Twitter for several years, and we do not face problems with composition changes. Consequently, we were able to conduct our analysis based on a good approximation of the content that circulated on Twitter. The initial step in our data collection involved retrieving the Tweets posted by each user within the Siegel et al. (2021) sample. In totality, our dataset encompassed 399 Million Tweets spanning from 2014 to the commencement of 2022, along with corresponding user profile information. We provide some summary statistics in Table A.1
@@
We developed a political Tweet classifier by fine-tuning the BERTweet model on a human-labeled dataset of political and non-political Tweets. To create the training dataset, we first compiled a list of political keywords and accounts associated with political figures. This produced a subset of data with a higher density of political Tweets, though many non-political Tweets remained. We then randomly sampled 10,000 Tweets and distributed them among five undergraduate research assistants for manual labeling. Coders were provided with labeling criteria and instructed to categorize each Tweet in a binary true/false format. Each subset included 200 shared Tweets to evaluate inter-coder reliability. We evaluated agreement using Cohen’s kappa coefficient, which ranges from 0 to 1, where 1 indicates perfect agreement and values near zero suggest agreement no better than chance. The average Cohen’s kappa score across all coder pairs was 0.76. We then use this dataset to train and evaluate our political Tweet classifier.
@@
To gauge the toxicity of these Tweets, we use Google’s Perspective API, a tool widely acknowledged for its efficacy in identifying hate speech (Wulczyn et al., 2017; Dixon et al., 2018). This state-of-the-art API assigns a toxicity score ranging from 0 to 1 across six distinct dimensions: general toxicity, severe toxicity, identity attacks, insults, profanity, and threats. The scores can be approximately interpreted as the probability that a randomly chosen user will classify content as toxic. For example, a score of 0.8 means that around 80% of users would judge the content to be toxic.
The Perspective API performs well in classifying toxic text and can assess Tweets in several languages, including English, Spanish, French, German, Portuguese, Italian, and Russian. In our dataset, English is, unsurprisingly, the predominant language, and we restrict our analysis to English-speaking Tweets. Instances where the API did not assign toxicity scores were primarily due to the absence of textual content, such as Tweets containing only hyperlinks. We provide examples of highly toxic Tweets in Table 1. The examples indicate that the Perspective API accurately identifies toxic content. As described in the main paper, we additionally use the Detoxify (Hanu and Unitary team, 2020) package and OpenAI’s Moderation API (OpenAI, 2024) as an alternative toxicity classification algorithm. Overall, we also find that the different models broadly agree in their toxicity evaluation of Tweets. The correlation between toxicity scores from the Perspective API and Detoxify is 0.85, while the correlation between the Perspective API and the Moderation API is 0.72.181818As the Moderation API does not report a direct Toxicity score, we always consider the maximum of all provided scores.
@@ B.1. Additional Details on Embeddings
To distill information from the raw text of Tweets into quantifiable data, we use three pre-trained models: BERTweet (Nguyen et al., 2020), RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), and DistilBert-Base-Multilingual-Cased-V2 (Reimers and Gurevych, 2019). These models all use on the BERT-style Transformer architecture, which has become a staple in text classification and natural language understanding tasks. The strength of these models comes from their utilization of an ”attention” mechanism, allowing them to evaluate words in the context of their surrounding text, thereby capturing the subtleties of language that are often lost in traditional analysis. We transform the raw Tweets into analyzable embeddings by tokenizing the text into its constituent word tokens. These tokens then serve as input to our models, which produce numerical representations—or embeddings—of each Tweet. These embeddings convey the semantic and syntactic nuances of the language used by Twitter users and form the backbone of our computational analysis. Further details on the specific attributes of these models within our study are provided in the following.
@@
The third model we are using in our analysis is DeBERTa (He et al., 2021). DeBERTa (Decoding-enhanced BERT with disentangled attention) introduced innovative mechanisms that refine the workings of BERT and RoBERTa models. Its distinctive feature lies in the disentangled attention mechanism, which considers the content and the position of words separately, offering a more nuanced understanding of the text. Each word is represented by dual vectors that capture what the word is and where it stands in a sentence. This allows for a better interpretation of language nuances. Furthermore, DeBERTa’s enhanced mask decoder predicts masked tokens by using their absolute positions, an improvement that aids in the pre-training process. With these advancements, DeBERTa improved the model’s pre-training efficiency and improved the performance across a variety of downstream tasks, like the generation of new text that mimics human speech.
DistilBert-Base-Multilingual-Cased-V2
DistilBERT-Base-Multilingual-Cased-V2 (Reimers and Gurevych, 2019) is a lighter and faster variant of BERT, specifically trained to handle multiple languages in a single model. DistilBERT applies knowledge distillation during training. This process reduces the model’s size and increases inference speed. The multilingual-cased version was trained on multiple languages using cased text, ensuring that it preserves distinctions in capitalization, which can be semantically relevant in many languages. As a result, this model is particularly well-suited for cross-lingual applications.
@@ Appendix C Additional Results
The following section describes additional results and robustness checks. We begin by showing that the average similarity of Tweets is not affected by content moderation. Further, we repeated our main analysis using alternative 1) embedding models, 2) toxicity scores, 3) engagement weighting, and 4) samples. We also demonstrate that our results are not driven by Tweets belonging to the “Insult” topic as created by the Top2Vec topic model. Lastly, we provide evidence from an alternative method to account for the toxicity dimension of the embedding space.
Removal of Toxic Content and Cosine Similarity
In Figure C.1, we report the average cosine similarity of Tweets after sequentially removing those above varying toxicity thresholds. Given the very high computational requirements of the similarity calculation, we conduct this analysis for a random subset of 500,000 Tweets. The results show that average similarity remains largely unaffected by the exclusion of toxic content. In unreported analyses, we confirm that this pattern holds when we focus on a subset of Tweets with the highest and lowest pairwise similarity.
Figure C.1: Removal of Toxic Content and Average Cosine Similarity
Notes: The figure shows the average cosine similarity of Tweets after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis.
@@
As a first robustness check, we reproduce our findings using the alternative embeddings described in Section B.1. This test rules out that our findings are driven by the particularities of the specific transformer model we have chosen, even though BERTweet is one of the standard choices for the analysis of English-speaking Twitter data. For this robustness exercise, we created new embeddings based on the RoBERTa, DeBERTa, and DistilBert-Base-Multilingual-Cased-V2 models and reconstructed the BCD based on these embeddings. The findings in Figure C.2 highlight that the findings are remarkably similar not only with regards to the overall patterns but also the magnitudes of the content-moderation-induced increases in the BCD.
@@
Notes: The figure shows the BCD computed using embeddings generated by DeBerta and RoBerta, and DistilBert-Base-Multilingual-Cased-V2 after the exclusion of toxic and random Tweets from the data.
@@
Engagement Weigting of Content
As a third robustness test, we weight the Tweets in our data by their engagement as measured by the number of Retweets when calculating the BCD. This allows us to arguably better account for for the frequency with which users would encounter the toxic tweets. The results from this analysis are shown in Figure C.4. We find that weighting Tweets by their engagement has no bearing on our results and leads to very similar magnitudes for the BCD.
Figure C.4: Robustness: Engagement Weighting
Notes: The figure shows the BCD after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis. We weight Tweets by the number of retweets when calculating the BCD. The blue line illustrates the BCD when an equivalent number of Tweets is excluded from the dataset at random.
@@
As a last robustness check, we repeat our main analysis based on three alternative samples of Tweets. First, we use 1 million randomly drawn Tweets from our English data without filtering for political content. Second, we use 1 million randomly drawn Tweets from a sample of politically interested German Twitter users. Third, we use 1 million randomly drawn Tweets from a sample of politically interested Italian Twitter users. To create embeddings for the German and Italian samples, we use the “bert-base-german-uncased” and “twitter-xlm-roberta-base” models, respectively. Toxicity in all cases was coded using the Perspectives API. For each of these samples, we repeat the analysis from Figure 1. The results are presented in Figure C.5. We find that independent sample content moderation appears to introduce distortions to the semantic space as measured by the BCD.
@@
The figure below reproduces Figure 1 when we remove all Tweets belonging to the “Insult” topic in advance. The results are virtually identical, suggesting that content moderation also introduces distortions of the semantic space beyond simply removing uncivil content.
@@
Alternative Method to Account for the Toxicity Dimension of the Embeddings
Our alternative approach for the orthogonalization of the embeddings with respect to the toxicity scores uses linear regressions of the following form:
$x_{d}=\alpha+\mathbf{Tox^{\prime}\beta}+\epsilon_{d}$
where $x_{d}\in\mathbf{X}$ is one of the $D$ embedding dimensions. $\mathbf{Tox}$ is a matrix containing 1000 indicators for the permilles of the toxicity distribution. [^19] We chose a non-parametric transformation of the toxicity scores to flexibly account for any potential non-linearities. As we show in a robustness check, the findings are almost identical if we instead residualize linearly with regard to the toxicity score. We estimate these regressions for all embedding dimensions $d\in D$ and replace the embedding dimensions with the regression residuals. The resulting residualized matrix $\mathbf{\tilde{X}}$ is orthogonal to the toxicity scores.
We then repeat our previous analysis based on the residualized embedding matrix $\mathbf{\tilde{X}}$. Figure C.7 visualizes the two different approaches to account for a Tweet’s toxicity. More specifically, we use regressions and either residualize the embeddings using permilles or the linear toxicity score. We again find that removing the toxicity component from the embedding space has little impact on our findings. The overall patterns are close to our original results. The BCD sharply increases once toxic Tweets are removed from the data.
Figure C.7: Controlling for Toxicity